OpenStudy (anonymous):

Peter is giving marbles to some children at a carnival. He has 5 red marbles, 4 blue marbles, and 3 yellow marbles. If Peter selects a marble randomly without looking, what is the probability that he will give a blue marble to the first child and then a yellow marble to the second child? 4 over 12 plus 3 over 12 is equal to 7 over 12 4 over 12 plus 3 over 11 is equal to 80 over 132 4 over 12 multiplied by 3 over 12 is equal to 12 over 144 4 over 12 multiplied by 3 over 11 is equal to 12 over 132

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@cupcakerain

OpenStudy (anonymous):

@Tabby_012 @night.soul45

OpenStudy (anonymous):

i will fan and medal

OpenStudy (misssunshinexxoxo):

Which do you believe it is?

OpenStudy (anonymous):

c or b @misssunshinexxoxo

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@LynFran

OpenStudy (kropot72):

The initial number of marbles is 12, of which 4 are blue. Therefore the probability of selecting a blue marble first is 4/12. Having selected a blue marble, there is a total of 11 marbles remaining of which 3 are yellow. Therefore the probability of a yellow marble being selected second is 3/11. The required probability is therefore given by \[\large \frac{4}{12}\times\frac{3}{11}\]

OpenStudy (anonymous):

0.0909090909 @kropot72

OpenStudy (anonymous):

i got it thanks

OpenStudy (kropot72):

np.